OpenStudy (anonymous):

Simplify: (m3/3n)/(-m4 /9n2)

OpenStudy (anonymous):

(m3/3n)/(-m4 /9n2) =(m3/3n)*(-9n2/m4) =-3n/m

OpenStudy (anonymous):

wat about (6y2 -6/ 8y2 +8y)/(3y -3 / 4y2 + 4y) @arch1

OpenStudy (anonymous):

are you sure the term (3y-3) is correct? should it be (3y^2 3)

ganeshie8 (ganeshie8):

i think so.. it should be 3y^2-3

OpenStudy (anonymous):

(6y2 -6/ 8y2 +8y)/(3y -3 / 4y2 + 4y) =(6y2 -6/ 8y2 +8y)*(4y2+4y/3y-3) =[6(y^2-1)/8y(y+1)]/[4y(y+1)/3(y-1)] if u solve it further by puttin (y2-1)=(y+1)(y-1) ull get ans. 2(y+1)/2

ganeshie8 (ganeshie8):

good work @arch1

OpenStudy (anonymous):

2 wil get cancel ans is (y+1)

OpenStudy (anonymous):

ty @ganeshie8

OpenStudy (anonymous):

assuming original expression is correct , arch's right, good work

OpenStudy (anonymous):

ty @gitahimart